There are 3 carriages in a train.the mass of each is 100 kg and mass of engine is 650kg and mass of engine driver is 50kg.the total friction by the track is 7500N.this force can divide among carriages and engine according to mass.What is the frictional force only on engine ?
1
Expert's answer
2020-01-20T05:20:55-0500

The friction force


"F_f=\mu N=\mu mg"

Hence, the total friction by the track is


"7500=(3\times 100+650+50)\mu g=1000\mu g"

The frictional force only on engine is given by

"F=650\mu g=650\times \frac{7500}{1000}=4875\:\rm N"